# Representations of a quantum-deformed Lorentz algebra, Clebsch-Gordan map, and Fenchel-Nielsen representation of quantum complex flat connections at level-$\bm{k}$

### v1

@@ 4.1 Fock-Goncharov coordinates and quantization
Let us focus on the triangulation in FIG.1. The representation of (4.1) on $\mathcal{H}\simeq L^{2}(\mathbb{R})\otimes\mathbb{C}^{k}$ can be constructed as the following: Firstly, we relate $\bm{X},\bm{Y}$ to $\bm{Z}_{1},\bm{Z}_{5}$ by
$-\bm{X}+i\pi=\bm{Z}_{1},\qquad-2\bm{Y}+i\pi=\bm{Z}_{5},\quad\text{i.e.}\quad\bm{z}_{1}=-\bm{x}^{-1},\qquad\bm{z}_{5}=-\bm{y}^{-2}.$
where the representations of $\bm{x},\bm{y}$ are the same as the above.
$\bm{y}=\exp\left[\frac{2\pi i}{k}(-ib\bm{\mu}-\bm{m})\right],\qquad\bm{x}=\exp\left[\frac{2\pi i}{k}(-ib\bm{\nu}-\bm{n})\right].$
@@
We may express $\bm{Z}_{1},\cdots,\bm{Z}_{6}$ in terms of $\bm{X},\bm{Y},L_{1},\cdots,L_{4}$
$\bm{Z}_{1}$ $=-\bm{X}+i\pi,\qquad\bm{Z}_{2}=-2L_{3}+2\bm{Y}+i\pi,$
$\bm{Z}_{3}$ $=L_{1}-L_{2}+L_{3}+L_{4}+\bm{X}+i\pi,$
$\bm{Z}_{4}$ $=-L_{1}+L_{2}+L_{3}+L_{4}-2\bm{Y}+i\pi,\qquad\bm{Z}_{5}=-2\bm{Y}+i\pi,$
$\bm{Z}_{6}$ $=-L_{1}-L_{2}-L_{3}-L_{4}+2\bm{Y}+i\pi$ (4.11)
@@
$\bm{z}_{1}$ $=-\bm{x}^{-1},\qquad\bm{z}_{2}=-\lambda_{3}^{-2}\bm{y}^{2},\qquad\bm{z}_{3}=-\frac{\lambda_{1}\lambda_{3}\lambda_{4}}{\lambda_{2}}\bm{x},$ (4.12)
$\bm{z}_{4}$ $=-\frac{\lambda_{2}\lambda_{3}\lambda_{4}}{\lambda_{1}}\bm{y}^{-2},\qquad\bm{z}_{5}=-\bm{y}^{-2},\qquad\bm{z}_{6}=-\frac{1}{\lambda_{1}\lambda_{2}\lambda_{3}\lambda_{4}}\bm{y}^{2}.$ (4.13)
@@
The tilded operator can be defined similarly
$\tilde{\bm{z}}_{1}$ $=-\tilde{\bm{x}}^{-1},\qquad\tilde{\bm{z}}_{2}=-\tilde{\lambda}_{3}^{-2}\tilde{\bm{y}}^{2},\qquad\tilde{\bm{z}}_{3}=-\frac{\tilde{\lambda}_{1}\tilde{\lambda}_{3}\tilde{\lambda}_{4}}{\tilde{\lambda}_{2}}\tilde{\bm{x}},$ (4.15)
$\tilde{\bm{z}}_{4}$ $=-\frac{\tilde{\lambda}_{2}\tilde{\lambda}_{3}\tilde{\lambda}_{4}}{\tilde{\lambda}_{1}}\tilde{\bm{y}}^{-2},\qquad\tilde{\bm{z}}_{5}=-\tilde{\bm{y}}^{-2},\qquad\tilde{\bm{z}}_{6}=-\frac{1}{\tilde{\lambda}_{1}\tilde{\lambda}_{2}\tilde{\lambda}_{3}\tilde{\lambda}_{4}}\tilde{\bm{y}}^{2}.$ (4.16)
where $\tilde{x},\tilde{y}$ are represented by
$\tilde{\bm{y}}=\exp\left[\frac{2\pi i}{k}(-ib^{-1}\bm{\mu}+\bm{m})\right],\qquad\tilde{\bm{x}}=\exp\left[\frac{2\pi i}{k}(-ib^{-1}\bm{\nu}+\bm{n})\right].$ (4.17)
and $\tilde{\lambda}_{1},\cdots,\tilde{\lambda}_{6}\in\mathbb{C}^{\times}$ are also constants and label the representation. We set $\tilde{\lambda}_{a}=\lambda_{a}^{*}$ so that
@@
representing the star structure. The above representation on $\mathcal{H}$ of the quantum FG coordinates may be denoted by $\varrho_{\{\lambda,\tilde{\lambda}\}}$, which is labelled by $\{\lambda,\tilde{\lambda}\}\equiv\{\lambda_{a},\tilde{\lambda}_{a}\}_{a=1}^{4}$.
In the following, we mainly focus on the case that $k=2N$ is an even number and $N$ is odd. In this case, we may decompose $\mathcal{H}$ into orthogonal subspaces
$\mathcal{H}\simeq\mathcal{H}_{+}\oplus\mathcal{H}_{-}$ (4.19)
where ±$\mathcal{H}_{\pm}\simeq L^{2}(\mathbb{R})\otimes\mathbb{C}^{N}$ contains ϕ±±subscriptitalic-$\phi_{\pm}\in\mathcal{H}_{\pm}$ satisfying
$\phi_{\pm}(\mu,m+N)=\pm\phi_{\pm}(\mu,m).$ (4.20)
$\mathcal{H}_{\pm}$ is the eigenspace of $e^{\pi i\bm{n}}$ with eigenvalue ±$\pm 1$. The quantization of FG coordinates relates to $\bm{x},\bm{y}^{2},\tilde{\bm{x}},\tilde{\bm{y}}^{2}$, which act irreducibly on each of $\mathcal{H}_{+}$ and $\mathcal{H}_{-}$. The full $\mathcal{H}$ is not irreducible with respect to the quantization of FG coordinates, although it is irreducible with respect to $\bm{x},\bm{y},\tilde{\bm{x}},\tilde{\bm{y}}$. Note that the action of $\bm{y}$ or $\tilde{\bm{y}}$ maps from $\mathcal{H}_{+}$ to $\mathcal{H}_{-}$ and vice versa.
@@ 4.2 S-cycle trace operator
We define the S-cycle to be the loop enclosing the 1st and 2nd holes (see FIG. 1). Classically, the trace of holonomy along the S-cycle can be written as a function of FG coordinates
$L_{s}$ $=$ $e^{-\frac{Z_{1}}{2}-\frac{Z_{3}}{2}-\frac{Z_{4}}{2}-\frac{Z_{5}}{2}}-e^{-\frac{Z_{1}}{2}+\frac{Z_{3}}{2}-\frac{Z_{4}}{2}-\frac{Z_{5}}{2}}+e^{-\frac{Z_{1}}{2}+\frac{Z_{4}}{2}+\frac{Z_{3}}{2}-\frac{Z_{5}}{2}}+e^{-\frac{Z_{1}}{2}+\frac{Z_{5}}{2}+\frac{Z_{3}}{2}+\frac{Z_{4}}{2}}$ (4.21)
$-\,e^{-\frac{Z_{1}}{2}+\frac{Z_{4}}{2}+\frac{Z_{5}}{2}+\frac{Z_{3}}{2}}+e^{\frac{Z_{1}}{2}+\frac{Z_{3}}{2}+\frac{Z_{4}}{2}+\frac{Z_{5}}{2}}.$
This is derived by using the ”snake rule” outlined in Appendix B. In the language of Teichmüeller theory, $L_{s}$ is the complexification of the FN length of the S-cycle, along which the 4-holed sphere is cut into two pairs of pants.
The quantization of the S-cycle trace by $Z_{a}\to\bm{Z}_{a}$ and inserting (4.1) leads to the operator
$\bm{L}_{s}$ $=$ $e^{-2\bm{Y}_{s}}+e^{2\bm{Y}_{s}}+e^{L_{1}-L_{2}+L_{3}+L_{4}+2\bm{Y}_{s}+\bm{X}}+e^{L_{1}-L_{2}+\bm{X}}+e^{L_{3}+L_{4}+\bm{X}}+e^{-2\bm{Y}_{s}+\bm{X}},$ (4.22)
$=$ $e^{-2\bm{Y}_{s}}+e^{2\bm{Y}_{s}}+e^{L_{3}+L_{4}}e^{-2\bm{Y}}{\bf q}^{-1}\left(e^{L_{1}-L_{2}-L_{3}-L_{4}}{\bf q}e^{2\bm{Y}}+1\right)\left(1+e^{2\bm{Y}}{\bf q}\right)e^{\bm{X}},$
where $2\bm{Y}_{s}=2\bm{Y}-L_{3}-L_{4}$. Its tilded counter-part $\tilde{\bm{L}}_{s}=\bm{L}_{s}^{\dagger}$ is obtained analogously
$\tilde{\bm{L}}_{s}$ $=$ $e^{-2\tilde{\bm{Y}}_{s}}+e^{2\tilde{\bm{Y}}_{s}}+e^{\tilde{L}_{3}+\tilde{L}_{4}}e^{-2\tilde{\bm{Y}}}{\tilde{\mathbf{q}}}^{-1}\left(e^{\tilde{L}_{1}-\tilde{L}_{2}-\tilde{L}_{3}-\tilde{L}_{4}}{\tilde{\mathbf{q}}}e^{2\tilde{\bm{Y}}}+1\right)\left(1+e^{2\tilde{\bm{Y}}}{\tilde{\mathbf{q}}}\right)e^{\tilde{\bm{X}}}.$ (4.23)
@@
$\bm{U}^{\prime}_{s}=\varphi\left(y^{2}e^{L_{1}-L_{2}-L_{3}-L_{4}},\tilde{y}^{2}e^{\tilde{L}_{1}-\tilde{L}_{2}-\tilde{L}_{3}-\tilde{L}_{4}}\right)\varphi\left(y^{2},\tilde{y}^{2}\right).$ (4.24)
@@
$\bm{U}^{\prime}_{s}\left[e^{L_{3}+L_{4}}e^{-2\bm{Y}}{\bf q}^{-1}\left(e^{L_{1}-L_{2}-L_{3}-L_{4}}{\bf q}e^{2\bm{Y}}+1\right)\left(1+e^{2\bm{Y}}{\bf q}\right)e^{\bm{X}}\right]\bm{U}_{s}^{\prime-1}=e^{-2\bm{Y}_{s}+\bm{X}}.$ (4.25)
@@
$\left(\begin{array}[]{c}Y\\ X\end{array}\right)\to\begin{pmatrix}1&0\\ -2&1\end{pmatrix}\left(\begin{array}[]{c}Y\\ X\end{array}\right),\qquad\bm{T}=e^{\frac{2\pi i}{k}\left(\bm{\mu}^{2}-\bm{m}^{2}\right)}.$ (4.33)
The unitary operator $\bm{T}$ satisfies
$\bm{T}e^{\bm{X}-2\bm{Y}}\bm{T}^{-1}=e^{\bm{X}}.$ (4.34)
@@
$\bm{L}_{s}^{\prime}=\bm{U}_{s}\bm{L}_{s}\bm{U}_{s}^{-1}=e^{-2\bm{Y}_{s}}+e^{2\bm{Y}_{s}}+e^{\bm{X}_{s}},\qquad\bm{U}_{s}=\bm{U}_{s}^{\prime}\bm{T},$ (4.35)
where $\bm{X}_{s}=\bm{X}+L_{3}+L_{4}$. A similar computation shows that $\bm{U}_{s}$ also simplify $\tilde{\bm{L}}_{s}$:
$\tilde{\bm{L}}_{s}^{\prime}=\bm{U}_{s}\tilde{\bm{L}}_{s}\bm{U}_{s}^{-1}=e^{-2\tilde{\bm{Y}}_{s}}+e^{2\tilde{\bm{Y}}_{s}}+e^{\tilde{\bm{X}}_{s}}.$ (4.36)
$\bm{L}_{s},\tilde{\bm{L}}_{s}$ can relate to the Casimir operator $\mathcal{Q}_{1}^{\prime\prime},\tilde{\mathcal{Q}}_{1}^{\prime\prime}$ in the representation theory of quantum group (recall (3.25) and (3.26)) by unitary transformations. Indeed, we further make a Fourier transform
[…]
such that
$\mathcal{F}^{-1}e^{\bm{Y}}\mathcal{F}=e^{\bm{X}},\qquad\mathcal{F}^{-1}e^{\bm{X}}\mathcal{F}=e^{-\bm{Y}}.$ (4.37)
Then
$\mathcal{F}^{-1}\bm{L}_{s}^{\prime}\mathcal{F}=e^{-2{\bm{X}}+L_{3}+L_{4}}+e^{2{\bm{X}}-L_{3}-L_{4}}+e^{-{\bm{Y}}+L_{3}+L_{4}}.$ (4.38)

### v2

@@ 4.1 Fock-Goncharov coordinates and quantization
Let us focus on the triangulation in FIG.1. The representation of (4.1) on $\mathcal{H}\simeq L^{2}(\mathbb{R})\otimes\mathbb{C}^{N}$ can be constructed as the following: Firstly, we relate $\bm{X},\bm{Y}$ to $\bm{Z}_{1},\bm{Z}_{5}$ by
$-\bm{U}=\bm{Z}_{1},\qquad-\bm{Y}+i\pi=\bm{Z}_{5},\quad\text{i.e.}\quad\bm{z}_{1}=\bm{u}^{-1},\qquad\bm{z}_{5}=-\bm{y}^{-1}.$
where the representations of $\bm{u},\bm{y}$ are the same as the above in (2.18).
@@
We may express $\bm{Z}_{1},\cdots,\bm{Z}_{6}$ in terms of $\bm{U},\bm{Y},L_{1},\cdots,L_{4}$
$\bm{Z}_{1}$ $=-\bm{U},\qquad\bm{Z}_{2}=-2L_{3}+\bm{Y}+i\pi,$
$\bm{Z}_{3}$ $=L_{1}-L_{2}+L_{3}+L_{4}+\bm{U}+2i\pi,$
$\bm{Z}_{4}$ $=-L_{1}+L_{2}+L_{3}+L_{4}-\bm{Y}+i\pi,\qquad\bm{Z}_{5}=-\bm{Y}+i\pi,$
$\bm{Z}_{6}$ $=-L_{1}-L_{2}-L_{3}-L_{4}+\bm{Y}+i\pi$ (4.11)
@@
$\bm{z}_{1}$ $=\bm{u}^{-1},\qquad\bm{z}_{2}=-\lambda_{3}^{-2}\bm{y},\qquad\bm{z}_{3}=\frac{\lambda_{1}\lambda_{3}\lambda_{4}}{\lambda_{2}}\bm{u},$ (4.12)
$\bm{z}_{4}$ $=-\frac{\lambda_{2}\lambda_{3}\lambda_{4}}{\lambda_{1}}\bm{y}^{-1},\qquad\bm{z}_{5}=-\bm{y}^{-1},\qquad\bm{z}_{6}=-\frac{1}{\lambda_{1}\lambda_{2}\lambda_{3}\lambda_{4}}\bm{y}.$ (4.13)
@@
The tilded operators can be defined similarly
$\tilde{\bm{z}}_{1}$ $=\tilde{\bm{u}}^{-1},\qquad\tilde{\bm{z}}_{2}=-\tilde{\lambda}_{3}^{-2}\tilde{\bm{y}},\qquad\tilde{\bm{z}}_{3}=\frac{\tilde{\lambda}_{1}\tilde{\lambda}_{3}\tilde{\lambda}_{4}}{\tilde{\lambda}_{2}}\tilde{\bm{u}},$ (4.15)
$\tilde{\bm{z}}_{4}$ $=-\frac{\tilde{\lambda}_{2}\tilde{\lambda}_{3}\tilde{\lambda}_{4}}{\tilde{\lambda}_{1}}\tilde{\bm{y}}^{-1},\qquad\tilde{\bm{z}}_{5}=-\tilde{\bm{y}}^{-1},\qquad\tilde{\bm{z}}_{6}=-\frac{1}{\tilde{\lambda}_{1}\tilde{\lambda}_{2}\tilde{\lambda}_{3}\tilde{\lambda}_{4}}\tilde{\bm{y}}.$ (4.16)
where $\tilde{\bm{u}},\tilde{\bm{y}}$ are represented as in (2.19). $\tilde{\lambda}_{1},\cdots,\tilde{\lambda}_{6}\in\mathbb{C}^{\times}$ are also constants and label the representation. We set $\tilde{\lambda}_{a}=\lambda_{a}^{*}$ so that
@@
representing the star structure. The above representation on $\mathcal{H}$ of the quantum FG coordinates may be denoted by $\varrho_{\{\lambda,\tilde{\lambda}\}}$ labelled by $\{\lambda,\tilde{\lambda}\}\equiv\{\lambda_{a},\tilde{\lambda}_{a}\}_{a=1}^{4}$.
4.2 Trace operators and relation to skein quantization
The flat connection is an $\mathrm{SL}(2,\mathbb{C})$-representation of the fundamental group on 4-holed sphere modulo conjugations. The fundamental group on 4-holed sphere $\pi_{1}$ is generated by the loops $\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4}$, each traveling counter-clockwisely around a hole and satisfying $\gamma_{1}\gamma_{2}\gamma_{3}\gamma_{4}=1$. The flat connection defines the $\mathrm{SL}(2,\mathbb{C})$ holonomies $H(\gamma_{1}),\cdots,H(\gamma_{4})$ modulo gauge transformations. On the moduli space of flat connections, the holonomies are $\mathrm{SL}(2,\mathbb{C})$-valued functions of logarithmic FG coordinates $Z_{1},\cdots,Z_{6}$ by the snake rule, as shown in Appendix B. These functions depend on the choice of gauge. But the traces of the holonomies are gauge invariant:
$A_{1}$ $=$ $\mathrm{Tr}\left(H(\gamma_{1})\right)=e^{-\frac{Z_{4}}{2}-\frac{Z_{6}}{2}+i\pi}+e^{\frac{Z_{4}}{2}+\frac{Z_{6}}{2}-i\pi},$
$B_{1}$ $=$ $\mathrm{Tr}\left(H(\gamma_{2})\right)=e^{-\frac{Z_{1}}{2}-\frac{Z_{3}}{2}-\frac{Z_{5}}{2}-\frac{Z_{6}}{2}+2i\pi}+e^{\frac{Z_{1}}{2}+\frac{Z_{3}}{2}+\frac{Z_{5}}{2}+\frac{Z_{6}}{2}-2i\pi},$ (4.18)
$C_{1}$ $=$ $\mathrm{Tr}\left(H(\gamma_{3})\right)=e^{-\frac{Z_{2}}{2}-\frac{Z_{5}}{2}+i\pi}+e^{\frac{Z_{2}}{2}+\frac{Z_{5}}{2}-i\pi},$ (4.19)
$D_{1}$ $=$ $\mathrm{Tr}\left(H(\gamma_{4})\right)=e^{-\frac{Z_{1}}{2}-\frac{Z_{2}}{2}-\frac{Z_{3}}{2}-\frac{Z_{4}}{2}+2i\pi}+e^{\frac{Z_{1}}{2}+\frac{Z_{2}}{2}+\frac{Z_{3}}{2}+\frac{Z_{4}}{2}-2i\pi}.$ (4.20)
In order to obtain a complete set of coordinate functions of the moduli space, we need to consider the traces of the S-cycle, T-cycle, and U-cycle holonomies. S-cycle is the loop enclosing the 1st and 2nd holes, T-cycle encloses the 2nd and 3rd holes (see FIG. 1), and U-cycle encloses the 1st and 3rd holes.
$L_{s}=-\mathrm{Tr}\left(H(\gamma_{1})H(\gamma_{2})\right)$ $=$ $e^{-\frac{Z_{1}}{2}-\frac{Z_{3}}{2}-\frac{Z_{4}}{2}-\frac{Z_{5}}{2}}-e^{-\frac{Z_{1}}{2}+\frac{Z_{3}}{2}-\frac{Z_{4}}{2}-\frac{Z_{5}}{2}}+e^{-\frac{Z_{1}}{2}+\frac{Z_{3}}{2}+\frac{Z_{4}}{2}-\frac{Z_{5}}{2}}$ (4.21)
$+e^{-\frac{Z_{1}}{2}+\frac{Z_{3}}{2}+\frac{Z_{5}}{2}-\frac{Z_{4}}{2}}-e^{-\frac{Z_{1}}{2}+\frac{Z_{3}}{2}+\frac{Z_{4}}{2}+\frac{Z_{5}}{2}}+e^{\frac{Z_{1}}{2}+\frac{Z_{3}}{2}+\frac{Z_{4}}{2}+\frac{Z_{5}}{2}},$
$L_{t}=-\mathrm{Tr}\left(H(\gamma_{2})H(\gamma_{3})\right)$ $=$ $e^{-\frac{Z_{1}}{2}-\frac{Z_{2}}{2}-\frac{Z_{3}}{2}-\frac{Z_{6}}{2}}-e^{\frac{Z_{1}}{2}-\frac{Z_{2}}{2}-\frac{Z_{3}}{2}-\frac{Z_{6}}{2}}+e^{\frac{Z_{1}}{2}+\frac{Z_{2}}{2}-\frac{Z_{3}}{2}-\frac{Z_{6}}{2}}$ (4.22)
$+e^{\frac{Z_{1}}{2}+\frac{Z_{6}}{2}-\frac{Z_{2}}{2}-\frac{Z_{3}}{2}}-e^{\frac{Z_{1}}{2}+\frac{Z_{2}}{2}+\frac{Z_{6}}{2}-\frac{Z_{3}}{2}}+e^{\frac{Z_{1}}{2}+\frac{Z_{2}}{2}+\frac{Z_{3}}{2}+\frac{Z_{6}}{2}},$
$L_{u}=-\mathrm{Tr}\left(H(\gamma_{1})H(\gamma_{3})\right)$ $=$ $-e^{-Z_{1}-\frac{Z_{2}}{2}-\frac{Z_{4}}{2}-\frac{Z_{5}}{2}-\frac{Z_{6}}{2}}-e^{-\frac{Z_{2}}{2}+\frac{Z_{4}}{2}-\frac{Z_{5}}{2}-\frac{Z_{6}}{2}}+e^{-Z_{1}+\frac{Z_{4}}{2}-\frac{Z_{2}}{2}-\frac{Z_{5}}{2}-\frac{Z_{6}}{2}}$ (4.23)
$-e^{-\frac{Z_{2}}{2}+\frac{Z_{5}}{2}-\frac{Z_{4}}{2}-\frac{Z_{6}}{2}}+e^{-Z_{1}+\frac{Z_{5}}{2}-\frac{Z_{2}}{2}-\frac{Z_{4}}{2}-\frac{Z_{6}}{2}}+2e^{-\frac{Z_{2}}{2}+\frac{Z_{4}}{2}+\frac{Z_{5}}{2}-\frac{Z_{6}}{2}}$
$-e^{-Z_{1}+\frac{Z_{4}}{2}+\frac{Z_{5}}{2}-\frac{Z_{2}}{2}-\frac{Z_{6}}{2}}-e^{Z_{1}+\frac{Z_{4}}{2}+\frac{Z_{5}}{2}-\frac{Z_{2}}{2}-\frac{Z_{6}}{2}}-e^{\frac{Z_{2}}{2}+\frac{Z_{4}}{2}+\frac{Z_{5}}{2}-\frac{Z_{6}}{2}}$
$+e^{Z_{1}+\frac{Z_{2}}{2}+\frac{Z_{4}}{2}+\frac{Z_{5}}{2}-\frac{Z_{6}}{2}}-e^{-\frac{Z_{2}}{2}+\frac{Z_{4}}{2}+\frac{Z_{5}}{2}+\frac{Z_{6}}{2}}+e^{Z_{1}+\frac{Z_{4}}{2}+\frac{Z_{5}}{2}+\frac{Z_{6}}{2}-\frac{Z_{2}}{2}}$
$-e^{Z_{1}+\frac{Z_{2}}{2}+\frac{Z_{4}}{2}+\frac{Z_{5}}{2}+\frac{Z_{6}}{2}}$
They satisfy a constraint [17, 25]
$\mathcal{C}$ $=$ $L_{s}L_{t}L_{u}-L_{s}^{2}-L_{t}^{2}-L_{u}^{2}-A_{1}B_{1}C_{1}D_{1}-A_{1}^{2}-B_{1}^{2}-C_{1}^{2}-D_{1}^{2}$ (4.24)
$-(A_{1}B_{1}+C_{1}D_{1})L_{s}-(B_{1}C_{1}+D_{1}A_{1})L_{t}-(C_{1}A_{1}+B_{1}D_{1})L_{u}+4=0.$
The polynomial ring of $A_{1},B_{1},C_{1},D_{1},L_{s},L_{t},L_{u}$ quotient by $\mathcal{C}$ gives the algebra of functions on the moduli space of $\mathrm{SL}(2,\mathbb{C})$ flat connections.
The quantization is given by applying (4.1) to $A_{1},B_{1},C_{1},D_{1},L_{s},L_{t},L_{u}$ and defines the following operators on $\mathcal{H}$
$\bm{A}_{1}$ $=$ $\lambda_{1}+\lambda_{1}^{-1},\qquad\bm{B}_{1}=\lambda_{2}+\lambda_{2}^{-1},\qquad\bm{C}_{1}=\lambda_{3}+\lambda_{3}^{-1},\qquad\bm{D}_{1}=\lambda_{4}+\lambda_{4}^{-1}$
$\bm{L}_{s}$ $=$ $e^{-L_{3}-L_{4}+\bm{Y}}+e^{L_{3}+L_{4}-\bm{Y}}-e^{L_{1}-L_{2}+\bm{U}+\bm{Y}}-e^{L_{3}+L_{4}+\bm{U}-\bm{Y}}-e^{L_{1}-L_{2}+\bm{U}}-e^{L_{3}+L_{4}+\bm{U}},$
$\bm{L}_{t}$ $=$ $e^{-L_{2}-L_{3}+\bm{Y}}+e^{L_{2}+L_{3}-\bm{Y}}-e^{L_{2}+L_{3}-\bm{U}-\bm{Y}}-e^{-L_{1}-2L_{3}-L_{4}-\bm{U}+\bm{Y}}-e^{L_{2}-L_{3}-\bm{U}}-e^{-L_{1}-L_{4}-\bm{U}}$
$\bm{L}_{u}$ $=$ $-e^{-L_{1}+L_{3}-\bm{U}-\bm{Y}}-e^{L_{1}+L_{3}+\bm{U}-\bm{Y}}-e^{L_{2}+L_{4}-\bm{U}-\bm{Y}}-e^{L_{2}+2L_{3}+L_{4}-\bm{U}-2\bm{Y}}-e^{L_{2}+2L_{3}+L_{4}+\bm{U}-2\bm{Y}}$ (4.25)
$-e^{L_{2}+2L_{3}+L_{4}+\bm{U}-\bm{Y}}-e^{-L_{1}-L_{3}-\bm{U}}-e^{L_{1}+L_{3}+\bm{U}}+e^{-L_{1}+L_{3}-\bm{Y}}+e^{L_{1}+L_{3}-\bm{Y}}+e^{L_{2}+L_{4}-\bm{Y}}$
$+2e^{L_{2}+2L_{3}+L_{4}-2\bm{Y}}+e^{L_{2}+2L_{3}+L_{4}-\bm{Y}}$
where $\bm{A}_{1},\bm{B}_{1},\bm{C}_{1},\bm{D}_{1}$ are proportional to identity operator on $\mathcal{H}$. These operators satisfy the following commutation relations
$\mathbf{q}^{-1}\bm{L}_{s}\bm{L}_{t}-\mathbf{q}\bm{L}_{t}\bm{L}_{s}$ $=$ $(\mathbf{q}^{-2}-\mathbf{q}^{2})\bm{L}_{u}-(\mathbf{q}-\mathbf{q}^{-1})(\bm{A}_{1}\bm{C}_{1}+\bm{B}_{1}\bm{D}_{1}),$ (4.26)
$\mathbf{q}^{-1}\bm{L}_{t}\bm{L}_{u}-\mathbf{q}\bm{L}_{u}\bm{L}_{t}$ $=$ $(\mathbf{q}^{-2}-\mathbf{q}^{2})\bm{L}_{s}-(\mathbf{q}-\mathbf{q}^{-1})(\bm{A}_{1}\bm{B}_{1}+\bm{C}_{1}\bm{D}_{1}),$ (4.27)
$\mathbf{q}^{-1}\bm{L}_{u}\bm{L}_{s}-\mathbf{q}\bm{L}_{s}\bm{L}_{u}$ $=$ $(\mathbf{q}^{-2}-\mathbf{q}^{2})\bm{L}_{t}-(\mathbf{q}-\mathbf{q}^{-1})(\bm{B}_{1}\bm{C}_{1}+\bm{A}_{1}\bm{D}_{1}),$ (4.28)
and the quantized constraint
$\bm{\mathcal{C}}$ $=$ $\mathbf{q}^{-1}\bm{L}_{s}\bm{L}_{t}\bm{L}_{u}-\mathbf{q}^{-2}\bm{L}_{s}^{2}-\mathbf{q}^{2}\bm{L}_{t}^{2}-\mathbf{q}^{-2}\bm{L}_{u}^{2}$ (4.29)
$-\mathbf{q}^{-1}(\bm{A}_{1}\bm{B}_{1}+\bm{C}_{1}\bm{D}_{1})\bm{L}_{s}-\mathbf{q}(\bm{B}_{1}\bm{C}_{1}+\bm{A}_{1}\bm{D}_{1})\bm{L}_{t}-\mathbf{q}^{-1}(\bm{A}_{1}\bm{C}_{1}+\bm{B}_{1}\bm{D}_{1})\bm{L}_{u}$
$-\bm{A}_{1}\bm{B}_{1}\bm{C}_{1}\bm{D}_{1}-\bm{A}_{1}^{2}-\bm{B}_{1}^{2}-\bm{C}_{1}^{2}-\bm{D}_{1}^{2}+\left(\frac{\mathbf{q}^{2}-\mathbf{q}^{-2}}{\mathbf{q}-\mathbf{q}^{-1}}\right)^{2}=0.$
The relations (4.26) - (4.29) coincide with the operator algebra $\mathcal{A}_{0,4}^{2}$ from the skein quantization of $\mathrm{SL}(2,\mathbb{C})$ flat connections on 4-holed sphere [17]. Therefore, the operators $\bm{A}_{1},\bm{B}_{1},\bm{C}_{1},\bm{D}_{1},\bm{L}_{s},\bm{L}_{t},\bm{L}_{u}$ define the representation of $\mathcal{A}_{0,4}^{2}$ carried by $\mathcal{H}$.
The Hermitian conjugates of $\bm{A}_{1},\bm{B}_{1},\bm{C}_{1},\bm{D}_{1},\bm{L}_{s},\bm{L}_{t},\bm{L}_{u}$ are denoted by the tilded operators $\tilde{\bm{A}}_{1},\tilde{\bm{B}}_{1},\tilde{\bm{C}}_{1},\tilde{\bm{D}}_{1},\tilde{\bm{L}}_{s},\tilde{\bm{L}}_{t},\tilde{\bm{L}}_{u}$. Their relations are given by (4.26) - (4.26) with $\mathbf{q}\to\tilde{\mathbf{q}}$ and the operators replaced by the tilded operators.
@@ 4.3 S-cycle trace operator
The S-cycle trace operator can be written as
$\bm{L}_{s}$ $=$ $e^{-\bm{Y}_{s}}+e^{\bm{Y}_{s}}-e^{L_{3}+L_{4}}e^{-\bm{Y}}{\bf q}^{-1}\left(e^{L_{1}-L_{2}-L_{3}-L_{4}}{\bf q}e^{\bm{Y}}+1\right)\left(1+e^{\bm{Y}}{\bf q}\right)e^{\bm{U}},$ (4.30)
where $\bm{Y}_{s}=\bm{Y}-L_{3}-L_{4}$. Its tilded partner $\tilde{\bm{L}}_{s}=\bm{L}_{s}^{\dagger}$ is obtained analogously
$\tilde{\bm{L}}_{s}$ $=$ $e^{-\tilde{\bm{Y}}_{s}}+e^{\tilde{\bm{Y}}_{s}}-e^{\tilde{L}_{3}+\tilde{L}_{4}}e^{-2\tilde{\bm{Y}}}{\tilde{\mathbf{q}}}^{-1}\left(e^{\tilde{L}_{1}-\tilde{L}_{2}-\tilde{L}_{3}-\tilde{L}_{4}}{\tilde{\mathbf{q}}}e^{\tilde{\bm{Y}}}+1\right)\left(1+e^{\tilde{\bm{Y}}}{\tilde{\mathbf{q}}}\right)e^{\tilde{\bm{U}}}.$ (4.31)
@@
$\bm{U}^{\prime}_{s}=\varphi\left(\bm{y}e^{L_{1}-L_{2}-L_{3}-L_{4}},\tilde{\bm{y}}e^{\tilde{L}_{1}-\tilde{L}_{2}-\tilde{L}_{3}-\tilde{L}_{4}}\right)\varphi\left(\bm{y},\tilde{\bm{y}}\right).$ (4.32)
@@
$\bm{U}_{s}^{\prime}\left[e^{L_{3}+L_{4}}{\bf q}^{-1}e^{-\bm{Y}}\left({\bf q}e^{L_{1}-L_{2}-L_{3}-L_{4}+\bm{Y}}+1\right)\left({\bf q}e^{\bm{Y}}+1\right)e^{\bm{U}}\right]\bm{U}_{s}^{\prime-1}=e^{-\bm{Y}_{s}+\bm{U}}.$ (4.33)
@@
$\left(\begin{array}[]{c}Y\\ U\end{array}\right)\to\begin{pmatrix}1&0\\ -1&1\end{pmatrix}\left(\begin{array}[]{c}Y\\ U\end{array}\right),\qquad\bm{T}=(-1)^{\bm{m}^{2}}e^{\frac{\pi i}{N}\left(\bm{\mu}^{2}-\bm{m}^{2}\right)}.$ (4.39)
The unitary operator $\bm{T}$ is well-defined on $\mathcal{H}$ and satisfies
$\bm{T}e^{\bm{U}-\bm{Y}}\bm{T}^{-1}=-e^{\bm{U}}.$ (4.40)
@@
$\bm{L}_{s}^{\prime}=\bm{U}_{s}\bm{L}_{s}\bm{U}_{s}^{-1}=e^{-\bm{Y}+L_{3}+L_{4}}+e^{\bm{Y}-L_{3}-L_{4}}+e^{\bm{U}+L_{3}+L_{4}},\qquad\bm{U}_{s}=\bm{U}_{s}^{\prime}\bm{T},$ (4.41)
A similar computation shows that $\bm{U}_{s}$ also simplify $\tilde{\bm{L}}_{s}$:
$\tilde{\bm{L}}_{s}^{\prime}=\bm{U}_{s}\tilde{\bm{L}}_{s}\bm{U}_{s}^{-1}=e^{-\tilde{\bm{Y}}+\tilde{L}_{3}+\tilde{L}_{4}}+e^{\tilde{\bm{Y}}-\tilde{L}_{3}-\tilde{L}_{4}}+e^{\tilde{\bm{U}}+\tilde{L}_{3}+\tilde{L}_{4}}.$ (4.42)